# Understanding and Mitigating Errors of LLM-Generated RTL Code

### v1

@@ VI-A Classification criteria for the two main error categories
When given natural language design descriptions, we consider an LLM to have correctly understood the requirements if it generates accurate high-level pseudocode (natural language format); We conclude the LLM failed to comprehend the design requirements when its generated pseudocode exhibits behavioral inconsistencies with the specifications. Conversely, if the LLM violates RTL syntax rules when translating correct pseudocode into low-level RTL code, we attribute this to insufficient mastery of RTL programming knowledge.
@@
Notably, some larger designs may contain both error types simultaneously (Fig. 14), where certain code segments comply with behavioral descriptions but violate RTL conventions (e.g., done assignment), while others exhibit behavioral inconsistencies due to knowledge limitations (e.g., data_reg recording); such designs are classified into both categories.
@@ VI-B Subclassification Criteria for Misinterpretation of Design Specifications
Initially, regarding long-context scenarios, we observed that directly prompting LLMs to generate RTL code based on lengthy contexts often resulted in certain code segments failing to accurately reflect high-level requirements. However, by manually decomposing the long context—retaining only the key descriptive text relevant to the erroneous segments while omitting unnecessary details—the LLM regenerated RTL code that correctly matched the specifications. As shown in Fig 23, after manual decomposition of the long context, the LLM successfully generated the correct code. This demonstrates that redundancy in long contexts is a limiting factor in LLMs’ ability to generate accurate RTL code.
@@
Similarly, in multimodal input scenarios, as illustrated in Fig 24, when a KMAP table is directly provided as input, the LLM struggles to correctly interpret the included information. However, when the table’s content is converted into natural language, the LLM successfully generates functionally accurate RTL code. This reveals a bottleneck in the LLM’s ability to comprehend tabular or other non-textual modalities.
@@
As shown in Fig 25, ambiguous design descriptions required manual refinement by removing contradictions and supplementing key information. This clarification significantly improved RTL generation accuracy. Remaining errors primarily stemmed from the LLM’s limited circuit knowledge, causing RTL-description mismatches. Cases like 26 demonstrate that providing essential circuit knowledge in the context helps generate correct RTL, confirming this knowledge gap as a major limitation.
Since we manually classify errors in LLM-generated RTL code, and given the inherent randomness or potential hallucinations in LLM outputs, some inaccuracies and imperfections in classification are inevitable. Nevertheless, it is undeniable that the aforementioned factors represent key bottlenecks limiting the LLM’s ability to generate correct RTL. Major challenges and future directions for applying LLMs to hardware design include determining effective methods to infuse circuit knowledge into LLMs, improving their robustness in generating code from ambiguous descriptions, and enhancing their comprehension of multimodal information and long-context understanding in hardware circuit design.
@@ VII Conclusion
In this paper, we systematically analyzed and categorized error causes in LLM-generated RTL code, and find that most errors stem not from the reasoning capabilities of LLMs, but rather from a lack of RTL programming knowledge, insufficient understanding of circuit concepts, ambiguous design descriptions, or misinterpretation of complex multimodal inputs. To address these, we propose a set of error correction techniques including: 1) RAG-based knowledge error mitigation, 2) rule-based description refinement, 3) multimodal data conversion, and 4) simulation-guided iterative debugging. We integrate the proposed error correction mechanisms into a representative LLM-based RTL code generation framework, achieving 91.0% accuracy on the VerilogEval Benchmark. This demonstrates that based on a series of strategies such as knowledge retrieval, without requiring additional training, can significantly mitigate errors in LLM-generated RTL code.

### v2

@@ VI-A Classification criteria for the two main error categories
When given natural language design descriptions, we consider an LLM to have correctly understood the requirements if it generates accurate high-level pseudocode. We conclude the LLM failed to comprehend the design requirements when its generated pseudocode exhibits behavioral inconsistencies with the specifications. Conversely, if the LLM violates RTL syntax rules when translating correct pseudocode into low-level RTL code, we attribute this to insufficient mastery of RTL programming knowledge.
@@
Notably, some larger designs may contain both error types simultaneously (Fig. 13), such designs are classified into both categories.
@@ VI-B Subclassification Criteria for Misinterpretation of Design Specifications
In long-context scenarios, directly prompting LLMs to generate RTL code from lengthy descriptions often leads to inaccuracies in certain code segments. However, manually decomposing the context—keeping only key requirements and removing unnecessary details (Fig 22)—enabled the LLM to produce correct RTL code. This shows that redundancy in long contexts limits the accuracy of LLM-generated RTL code.
@@
Similarly, as shown in Fig. 23, when the content of a KMAP table is converted into natural language, the LLM successfully generates functionally accurate RTL code, revealing a bottleneck in its ability to directly comprehend tabular or other non-textual modalities.
@@
As shown in Fig. 24, ambiguous design descriptions required manual refinement by removing contradictions and supplementing key information. This clarification significantly improved RTL generation accuracy. Remaining errors primarily stemmed from the LLM’s limited circuit knowledge, causing RTL-description mismatches. Cases like Fig. 25 demonstrate that providing essential circuit knowledge in the context helps generate correct RTL, confirming this knowledge gap as a major limitation.
@@ VII Conclusion
In this paper, we systematically analyzed and categorized error causes in LLM-generated RTL code, and find that most errors stem not from the reasoning capabilities of LLMs, but rather from a lack of RTL programming knowledge, insufficient understanding of circuit concepts, ambiguous design descriptions, or misinterpretation of complex multimodal inputs. To address these, we propose a set of error correction techniques including: 1) rule-based description refinement, 2) multimodal data conversion, 3) RAG-based knowledge error mitigation, and 4) simulation-guided iterative debugging. We integrate the proposed error correction mechanisms into a representative LLM-based RTL code generation framework, achieving 98.1% accuracy on the VerilogEval Benchmark. This demonstrates that based on a series of strategies such as knowledge retrieval, without requiring additional training, can significantly mitigate errors in LLM-generated RTL code.